### v2

# WaterSeeker: Efficient Detection of Watermarked Segments in Large Documents

Abstract
Watermarking algorithms for large language models (LLMs) have attained high accuracy in detecting LLM-generated text. However, existing methods primarily focus on distinguishing fully watermarked text from non-watermarked text, overlooking real-world scenarios where LLMs generate only small sections within large documents. In this scenario, balancing time complexity and detection performance poses significant challenges. This paper presents WaterSeeker, a novel approach to efficiently detect and locate watermarked segments amid extensive natural text. It first applies an efficient anomaly extraction method to preliminarily locate suspicious watermarked regions. Following this, it conducts a local traversal and performs full-text detection for more precise verification. Theoretical analysis and experimental results demonstrate that WaterSeeker achieves a superior balance between detection accuracy and computational efficiency. Moreover, WaterSeeker’s localization ability supports the development of interpretable AI detection systems. This work pioneers a new direction in watermarked segment detection, facilitating more reliable AI-generated content identification. Our code is available at https://github.com/THU-BPM/WaterSeeker.
1 Introduction
As the quality of text generated by large language models (LLMs) improves, these models address various practical challenges while also raising concerns such as misinformation (Liu et al., 2023b; Chen & Shu, 2024) and copyright infringement (Rillig et al., 2023). To tackle these issues, LLM watermarking technology has emerged, embedding specific information (watermarks) during text generation for accurate detection through specialized algorithms. Existing watermarking methods mainly focus on distinguishing fully watermarked text from non-watermarked content (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b; Lu et al., 2024; Aaronson & Kirchner, 2022; Kuditipudi et al., 2024). However, in real-world scenarios, LLMs often generate only brief segments within longer documents, making it essential to detect these watermarked sections among large amounts of non-watermarked text.
Most previous algorithms relied on computing statistics across the entire document for detection, which we refer to as full-text detection methods. For instance, some approaches categorize vocabulary into green and red lists, applying a bias towards the selection of green tokens during generation and using metrics like the z-score to measure their proportion throughout the entire text (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b). Other methods utilize pseudo-random numbers to guide token sampling, calculating correlations across the entire document during detection (Aaronson & Kirchner, 2022; Christ et al., 2024; Kuditipudi et al., 2024). Ultimately, the presence of a watermark is determined by whether these statistics exceed a specified threshold.
However, these full-text detection methods struggle when small amounts of watermarked text are mixed with large volumes of natural text due to dilution effects, as illustrated in Figure 1(a). To the best of our knowledge, the only existing work addressing this scenario is the WinMax (Kirchenbauer et al., 2024) algorithm, which examines all possible window sizes and traverses the full text for each size. However, this method suffers from high time complexity. Additionally, our self-constructed baseline, the Fixed-Length Sliding Window (FLSW) method, faces the issue of watermark strength dilution due to inflexible window sizes, which negatively impacts detection performance.
(a) Full-text detection methods often struggle in watermarked segment detection scenarios due to the dilution effect.
(b) An illustration of the WaterSeeker algorithm, using “first locate, then detect” strategy to balance time complexity and detection performance.
Figure 1: An overview of watermark detection scenarios and the proposed WaterSeeker algorithm.
To address these issues, we propose a novel watermark detection method named WaterSeeker, which is adaptable to various watermarking techniques. This method achieves a superior balance between time complexity and detection performance when identifying watermarked segments within large documents. The core concept of WaterSeeker is “first locate, then detect”. As illustrated in Figure 1(b), it initially employs a low-complexity anomaly points extraction algorithm to identify suspected watermark regions. This process narrows the detection target from a long text down to a small segment that encompasses the ground truth segment. Next, a local traversal is performed on the localization result. Within each window, full-text watermark detection is conducted, and the highest confidence result is compared to a threshold for the final determination. Theoretical analysis indicates that this coarse-to-fine process has the potential to achieve optimal detection performance while maintaining the lowest possible complexity for solving this problem.
In the experiment, we compared the effectiveness of WaterSeeker and the baselines in detecting watermarked segments of random lengths inserted into large documents, as well as their time complexity. The results show that WaterSeeker significantly surpasses the baseline methods in effectively balancing time complexity and detection performance. Additionally, more detailed analysis includes an adaptability assessment of WaterSeeker under varying watermark strengths and segment lengths, as well as an evaluation of its two-stage utility via ablation study. Furthermore, we discuss WaterSeeker’s contribution to building more interpretable AI detection systems.
In summary, the contributions of this work are as follows:
  • We comprehensively define the new scenario of detecting watermarked segments in large documents, specifying the inputs and outputs of the algorithms designed to address this problem, as well as the evaluation metrics and the construction process for the test dataset.
  • We propose WaterSeeker, a novel and general watermark detection method that effectively identifies watermarked segments in large documents, addressing the challenges posed by dilution effects in non-watermarked text.
  • WaterSeeker achieves a superior balance between time complexity and detection effectiveness, with strong adaptability to varying watermark strengths and lengths, significantly outperforming existing baselines.
  • WaterSeeker not only provides accurate detection results but also offers precise localization outcomes, contributing to the development of more interpretable AI detection systems.
2 Related Work
Watermarking for large language models (LLMs) embeds subtle, algorithmically detectable signals in outputs to identify LLM-generated text, making it crucial for mitigating potential misuse of LLMs. Currently, mainstream watermarking methods involve intervening in the inference phase of LLMs by modifying the logits output or influencing the token sampling process to produce watermarked text (Liu et al., 2023a; Pan et al., 2024; Liu et al., 2024c).
The KGW method (Kirchenbauer et al., 2023) modifies logits by categorizing the vocabulary into green and red lists, applying a bias towards green tokens during text generation. The bias value is typically determined by the parameter $\delta$, which reflects the watermark strength. For detection, it first labels each token as red or green and calculates the z-score of the entire text to reflect the proportion of green tokens; if this proportion exceeds a specified threshold, the text is classified as watermarked. In subsequent work, numerous enhancements have been suggested to improve text quality (Hu et al., 2024; Wu et al., 2023; Takezawa et al., 2023), increase information capacity (Wang et al., 2024; Yoo et al., 2023; Fernandez et al., 2023), boost robustness against watermark removal attacks (Zhao et al., 2024; Liu et al., 2024b; Ren et al., 2024; He et al., 2024), adapt to low-entropy conditions (Lee et al., 2023; Lu et al., 2024), and facilitate public detection (Liu et al., 2024a; Fairoze et al., 2023).
On the other hand, the Aar method (Aaronson & Kirchner, 2022) proposes utilizing pseudo-random sequences to guide the token sampling process. During each token sampling step, a pseudo-random vector $u\sim\text{Uniform}([0,1])^{|V|}$ is generated based on the previously sampled tokens. Assuming that the probability vector provided by the LLM is $p$, the method performs exponential sampling to select the token $i$ that maximizes $u_{i}^{1/p_{i}}$. Here, the watermark strength is reflected by a sampling temperature, which affects the strictness of the exponential sampling. For detection, it also employs global statistics: it sums the correlation values of each token with the pseudo-random vector and then performs gamma transformation to derive the detection confidence. Building on this, Kuditipudi et al. (2024) introduced the use of edit distance to calculate correlation, while Hou et al. (2024a) and Hou et al. (2024b) incorporated contextual semantic information to guide the sampling process, thereby enhancing detection robustness.
Despite the high accuracy of watermarking algorithms for detecting LLM-generated text, previous studies have primarily focused on distinguishing between fully watermarked and non-watermarked text, overlooking the possibility that LLMs may only generate small segments within large documents. In such cases, watermark detection algorithms based on full-text statistics can fail due to the dilution effect. A few studies have mentioned copy-paste attack (Kirchenbauer et al., 2024; Yoo et al., 2023; Wang et al., 2024), where a certain proportion of watermarked text is mixed into non-watermarked content, resembling our scenario. Yoo et al. (2023) and Wang et al. (2024) tested the robustness of their methods against copy-paste attacks by mixing 10% to 50% watermarked text into non-watermarked text. However, since they did not specifically design detection mechanisms for this scenario, their results indicated a lack of robustness against such attack.
In existing studies, WinMax (Kirchenbauer et al., 2024) is the only watermark detection algorithm specifically designed for this scenario. It examines all possible window sizes and traverses the entire text for each size to identify the maximum local score for threshold comparison. While this method shows some effectiveness, its high time complexity renders it impractical for real-world applications. Therefore, to achieve effective detection of watermarked segments inserted into large documents while maintaining controllable time complexity, we propose WaterSeeker, a novel and general method that uses “first locate then detect” strategy. We validate its superiority through comparisons with full-text detection, WinMax, and a self-constructed fixed-length sliding window detection method.
3 Problem Formulation
3.1 Definition
The problem of detecting watermarked segments in a long document is defined as follows: Let $N$ be a long text with a randomly inserted watermarked segment of length $L$ (where $L$ is a random variable). We denote the starting index of the watermarked segment in $N$ as $s$ and the ending index as $e$, such that: $L=e-s+1$. The objective is to determine the presence and location of a watermarked segment in $N$. This can be framed as a binary classification problem with additional localization, where the detection algorithm outputs:
[…]
Here, ‘has_watermark’ is a boolean value indicating the presence of a watermark, and ‘indices’ is a list of pairs $(s^{\prime},e^{\prime})$ representing the start and end indices of detected watermarked segments.
3.2 Evaluation
A watermark is considered successfully detected if:
[…]
  • There exists at least one detected segment $(s^{\prime},e^{\prime})$ such that the Intersection over Union (IoU) with the ground truth segment $(s,e)$ exceeds a specified threshold $\theta$:
    $\text{IoU}=\cfrac{L_{\text{intersection}}}{L_{\text{union}}}>\theta.$ (1)
Based on this evaluation, we will report the following metrics of the binary classification result: False Positive Rate (FPR), False Negative Rate (FNR), and F1 Score. Additionally, we will include the average IoU between the detected segments and ground truth segments as a supplementary metric, further demonstrating the accuracy of the localization.
4 Baseline Methods
4.1 Full-text Detection
[…]
4.2 WinMax
WinMax (Kirchenbauer et al., 2024) is a watermark detection algorithm proposed on the basis of KGW to address copy-paste attack. It involves iterating through all possible window sizes, and for each window size, the entire text is traversed to calculate the z-score for each local window, taking the maximum z-score and comparing it against a specified threshold. In practical applications, a minimum window size $W_{min}$ and a maximum window size $W_{max}$ are typically set to control the number of iterations during the traversal. The detection process can be described by the following formula:
$z_{\text{win-max}}=\max_{w\in[W_{min},W_{max}]}\left(\max_{i}\left(\frac{|s|_{G,i}-\gamma w}{\sqrt{\gamma(1-\gamma)w}}\right)\right),$ (2)
where $w$ is the length of the local window, and $|s|_{G,i}$ represents the count of green tokens within the $i$-th local window of length $w$. For other watermarking algorithms, such as Aar, a similar approach can also be implemented. The time complexity of the WinMax detection method is relatively high, costing $O((W_{max}-W_{min})\times N)$, where $N$ is the total length of the text. Due to the uncertain length of the watermarked segment inserted into the document, if there is a large difference between $W_{max}$ and $W_{min}$, the worst-case complexity can reach $O(N^{2})$.
4.3 Fix-Length Sliding Window
Fix-Length Sliding Window (FLSW) detection method is a self-constructed, straightforward approach to address the watermark detection problem in this scenario. It involves using a fixed-length window to traverse the entire text, calculating statistics within the local window. If the statistics within a local window exceed a specified threshold, the variable ‘has_watermarked’ is set to true. For localization, during the traversal, each pair of indices $(s,e)$ that meets the condition is recorded. Finally, a contiguous method with a certain tolerance is employed to provide the localization results. The pseudocode for the aforementioned baselines can be found in Appendix A.
Although the concept of FLSW is relatively simple and straightforward, the fixed-length nature makes it unable to adapt to varying watermark lengths. This limitation can result in dilution effects when the window size is either too large or too small. In Section 5.1, we will analyze this dilution effect in detail using KGW (Kirchenbauer et al., 2023) and Aar (Aaronson & Kirchner, 2022) as examples, providing theoretical support for WaterSeeker.
5 Proposed Method: WaterSeeker
5.1 Theoretical Foundation: Gold Index is the Best
In this section, we will conduct a theoretical analysis to demonstrate that using the actual start and end indices (the gold index) for watermark detection yields the highest expected detection rate, using KGW and Aar as examples. As mentioned in Section 2, most watermarking algorithms are slight modifications of these two methods, maintaining the fundamental principles of statistical calculation. Therefore, analyzing these two approaches provides a broad applicability.
For the KGW method, assuming $\gamma_{1}>\gamma$ is the proportion of green tokens in the watermarked part. Let’s analyze the effect of window size $W$ on this statistic:
  • When $W<L$ (window size smaller than watermark length), we have $E[|s|_{G}]=W\gamma_{1}$. Therefore,
    $E[z_{W}]=\frac{W\gamma_{1}-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\sqrt{W}\cdot\frac{\gamma_{1}-\gamma}{\sqrt{\gamma(1-\gamma)}},$ (3)
    where the expectation of z-score is proportional to $\sqrt{W}$.
  • When $W>L$ (window size larger than watermark length), we have $E[|s|_{G}]=L\gamma_{1}+(W-L)\gamma$. Therefore
    $E[z_{W}]=\frac{L\gamma_{1}+(W-L)\gamma-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\frac{L(\gamma_{1}-\gamma)}{\sqrt{\gamma(1-\gamma)W}},$ (4)
    where the expectation of z-score is proportional to $1/\sqrt{W}$, decreasing as $W$ increases.
From this, we can conclude that when $W=L$, the z-score reaches its maximum.
During detection, we aim for a higher z-score for positive cases while setting an appropriate threshold to balance the false positive rate. Next, we will analyze the constraints on the z-threshold $z^{*}$ when the false positive rate within the specified window is set to be lower than a target value $\alpha$.
We start with the binomial distribution $B(W,\gamma)$ that describes the number of green tokens in a window of size $W$. For large sample sizes (typically $W>50$), we can apply the Central Limit Theorem to approximate this binomial distribution with a normal distribution $N(W\gamma,W\gamma(1-\gamma))$. We then standardize this to a z-score, which follows the standard normal distribution $N(0,1)$. The false positive rate alpha represents the area in the right tail of this distribution beyond $z^{*}$. Mathematically, this is expressed as $\alpha=1-\Phi(z^{*})$, where $\Phi$ is the cumulative distribution function (CDF) of the standard normal distribution. Solving for $z^{*}$, we get:
$z^{*}=\Phi^{(-1)}(1-\alpha),$ (5)
which is a constant value for different $W$.
For the Aar method, the p-value is calculated by applying a gamma transformation to the sum of correlation values of all the tokens:
[…]
Due to the complexity of the Gamma-Transformation, the detailed proof can be found in Appendix B. Based on the above proof, we simulated the expected score of the statistic and the corresponding threshold as they vary with $W$ for $L=200$ using real data, as shown in Figure 2. It demonstrates that when the window size corresponds to the gold index, the highest detection rate can be achieved while maintaining an acceptable FPR. Therefore, the watermark dilution effect refers to the situation where the window used for detection is not optimal, resulting in the computed statistics representing watermark strength not reaching their peak or minimum, which in turn affects detection performance.
(a) KGW, $\alpha=10^{-6},\gamma=0.5,\gamma_{1}=0.75$
(b) Aar, $\alpha=10^{-6},\mu=1.0,\mu_{1}=1.6$
Figure 2: Expected score of the statistics and the corresponding threshold across varying $W$.
5.2 Suspicious Region Localization
Based on the theoretical analysis presented in Section 5.1, the design concept of WaterSeeker employs a coarse-to-fine process to gradually approximate the gold index, thereby achieving high detection performance. In the first step (coarse step), WaterSeeker uses a localization algorithm to identify suspicious watermarked regions. This process narrows the detection target to a small segment that encompasses the ground truth segment, while ensuring relatively small deviation from the gold index. Below is the detailed process of the localization algorithm:
  1. 1.
    Score List Computation: We utilize a small sliding window (i.e., $W=50$) to efficiently traverse the entire text. For each position of the window, we compute a statistical measure that reflects the average watermark intensity. For example, in the KGW method, this statistic is represented as $|s|_{G}/W$, indicating the density of green tokens inside the window. Meanwhile, in the Aar method, it is calculated as $\cfrac{\sum\log(1/(1-u))}{W}$, representing the average correlation value between tokens and the pseudo-random sequence within the window. This results in a score list $s$ of length $N-W+1$, where $s_{i}$ represents the average watermark intensity of $n_{i}$ to $n_{i}+W$.
  2. 2.
    Anomaly Extraction: In this step, we design an anomaly extraction algorithm to identify outliers from the score list. We calculate the mean of the score list, the top-k mean, and the standard deviation, denoted as $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$, respectively. The extracted outliers are the points that satisfy the following condition, where $\theta_{1}$ and $\theta_{2}$ are hyperparameters:
    $\text{score}>s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2}).$ (7)
    In samples containing a watermarked segment, the first term in the max expression becomes significant. Introducing the top-k mean allows for adaptive adjustment of the deviation from the mean based on different watermark strengths. Conversely, in non-watermarked samples, the difference between the top-k mean and the mean is relatively small. In this case, to reduce false positives, a point must deviate from the mean by at least $s_{\text{std}}\cdot\theta_{2}$ to be considered an anomaly.
  3. 3.
    Fragment Connection: In this step, we use a connecting method with a certain tolerance to link nearby outliers, then filter out segments that are too short, returning a list of indices.
The scores within watermarked segments tend to be relatively stable and close to the top-k mean. Concurrently, the scores corresponding to segments outside the watermark also remain stable and approximately equal to the overall mean (because a large volume of natural text dilutes the overall watermark strength). As a result of these characteristics, when we use $(s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1}\quad(\theta_{1}<1)$ as a criterion for distinction, the extracted abnormal segment’s start and end points ($s^{\prime}$ and $e^{\prime}$) generally satisfy the following conditions: $s^{\prime}\in(s-W,s)$ and $e^{\prime}\in(e,e+W)$. This ensures that the extracted suspicious watermarked regions are likely to cover the actual segments while keeping the starting and ending deviations within a window size.
Algorithm 1 WaterSeeker Algorithm
1:procedure SuspiciousRegionLocalization(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$)
2:     Compute score list using sliding window
3:     Calculate $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$
4:     Detect anomalies using threshold:
5:     $s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2})$
6:     Connect nearby outliers and filter short segments
7:     return filteredSegments
8:end procedure
9:procedure LocalTraverseDetection(tokens, suspiciousRegions, $W$, threshold)
10:     hasWatermark $\leftarrow$ False
11:     indices $\leftarrow$ [ ]
12:     for $(s^{\prime},e^{\prime})$ in suspiciousRegions do
13:         maxScore = $\infty$- ∞, maxIndice = ()
14:         for $s\in[s^{\prime},s^{\prime}+W)$ and $e\in(e^{\prime}-W,e^{\prime}]$ do
15:              if DetectWatermark(tokens[$s:e$])$>$ maxScore then
16:                  maxScore = DetectWatermark(tokens[$s:e$])
17:                  maxIndice = $(s,e)$
18:              end if
19:         end for
20:         if maxScore $>$ threshold then
21:              hasWatermark $\leftarrow$ True
22:              indices.append(maxIndice)
23:         end if
24:     end for
25:     return hasWatermark, indices
26:end procedure
27:procedure WaterSeeker(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$, threshold)
28:     suspiciousRegions $\leftarrow$ SuspiciousRegionLocalization(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$)
29:     hasWatermark, indices $\leftarrow$ LocalTraverseDetection(tokens, suspiciousRegions, W, threshold)
30:     return hasWatermark, indices
31:end procedure
5.3 Local Traverse Detection
After obtaining the localization results from the first step, the second phase (fine step), local traverse detection, conducts a more detailed verification. For each $(s^{\prime},e^{\prime})$ pair in the localization results, traverse inwards through the segments where the start point falls within $[s^{\prime},s^{\prime}+W)$ and the end point within $(e^{\prime}-W,e^{\prime}]$. Based on the previous analysis of the ranges of $s^{\prime}$ and $e^{\prime}$, this traversal has a high probability of reaching the gold index. Perform full-text detection sequentially on these segments, and select the statistic with the most significance (e.g., for KGW, choose the segment with the largest z-score, and for Aar, the segment with the smallest p-value) to compare with the threshold. The pseudocode for the entire WaterSeeker algorithm can be found in Algorithm 1.
5.4 Time Complexity Analysis
Time Complexity Analysis of WaterSeeker:
  1. 1.
    Suspicious Region Localization: $O(N)$. This step involves three sub-steps: score list computation, anomaly detection, and fragment connection. Each of these sub-steps has a time complexity of $O(N)$, where $N$ is the length of the text.
  2. 2.
    Local Traverse Detection: $O(W^{2})$. For each suspicious region identified, this step examines $W^{2}$ windows, where $W$ is the window size.
The total time complexity of WaterSeeker is $O(N+W^{2})$. In practice, $W^{2}$ is typically maintained at a value lower than $T$, as even a slightly larger window (i.e., $W=50$) is sufficient to provide a relatively smooth and low-noise representation of the surrounding watermark intensity. Therefore, the dominant term in most cases is $O(N)$, making the overall time complexity of WaterSeeker $O(N)$.
Lower Bound Complexity Analysis for the Problem: To detect watermarked segments in a long text, any algorithm must examine each token in the text at least once. This necessitates at least one full pass through the text, establishing a lower bound of $\Omega(N)$ for the time complexity of this problem.
In conclusion, the WaterSeeker algorithm, with its $O(N)$ complexity in typical cases, achieves a time complexity that matches the theoretical lower bound of the problem, indicating that it is asymptotically optimal for the task of detecting watermarked segments in long texts.
6 Experiment
6.1 Experiment Settings
Watermarking Methods and Language Models: We selected two representative watermarking algorithms, KGW (Kirchenbauer et al., 2023) and Aar (Aaronson & Kirchner, 2022), each with three strength levels. For KGW, the strength is determined by parameter $\delta$, with values 2.0, 1.5, and 1.0 (from strong to weak). For Aar, the strength is controlled by parameter temperature, with values 0.3, 0.2, and 0.1 (from strong to weak). We used Llama-2-7b (Touvron et al., 2023) and Mistral-7b (Jiang et al., 2024) as generation models.
Dataset Construction: Following previous studies, the C4 dataset (Raffel et al., 2020) was used for prompts, with the first 30 tokens of each entry serving as input. Watermarked segments of random length (100 to 400 tokens) were generated using randomly selected watermark strengths. For positive examples, one such segment was randomly inserted into each 10,000-token Wikipedia passage (Foundation, ). Negative examples consist of unmodified 10,000-token Wikipedia corpus. Based on this procedure, four datasets were created, each containing 300 positive and 300 negative examples: KGW-llama, KGW-mistral, Aar-llama, and Aar-mistral. These datasets employ different combinations of watermarking methods and LLMs. The specific distributions of watermark strengths and lengths are detailed in Appendix D.
It is important to note that watermarked segments with higher strength and longer length are inherently easier to detect and demonstrate better robustness against the dilution effect. Therefore, to showcase the superiority and adaptability of WaterSeeker, each dataset incorporates watermarked segments of varying strengths and lengths.
Baselines: As introduced in Section 4, Full-text Detection and WinMax (Kirchenbauer et al., 2024) are chosen, as well as Fix-Length Sliding Window method with $W$ of 100, 200, 300 and 400.
Hyper-parameters: The parameters related to WaterSeeker are as follows: $W=50,k=20$, with a tolerance for fragment connection set to 100. The parameter $\theta_{1}$ is set to 0.5, and $\theta_{2}$ is set to 1.5 for both algorithms. The threshold selection within the specified window is detailed in Appendix C. Notably, careful threshold selection is crucial for maintaining an acceptable false positive rate, as traversing long texts is prone to accumulating false positives. Evaluation-related parameters: $\theta$ mentioned in Section 3.2 is set to 0.5.
6.2 Detection Performance Analysis
Table 1: We compared the detection performance of WaterSeeker with other methods, including Plain Detect, WinMax (Kirchenbauer et al., 2024), and FLSW. The reported metrics include False Positive Rate (FPR), False Negative Rate (FNR), F1 score, and the average Intersection over Union (IoU) between detected segments and ground truth segments in positive samples.
Model Method KGW Aar
FPR FNR F1 IoU FPR FNR F1 IoU
Llama-2-7b Full-text Detection 0.000 1.000 [0.987] 0.000 [0.026] 0.000 0.000 1.000 [0.967] 0.000 [0.065] 0.001
WinMax 0.017 0.273 0.834 0.661 0.017 0.410 0.734 0.537
FLSW-100 0.003 0.473 0.688 0.448 0.003 0.547 0.622 0.363
FLSW-200 0.003 0.440 0.716 0.417 0.000 0.577 0.595 0.338
FLSW-300 0.007 0.683 0.479 0.313 0.000 0.783 0.356 0.265
FLSW-400 0.003 0.897 0.187 0.230 0.003 0.933 0.125 0.191
WaterSeeker(Ours) 0.017 0.303 0.813 0.629 0.010 0.433 0.719 0.510
Mistral-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.000 0.000 1.000 [0.973] 0.000 [0.052] 0.001
WinMax 0.010 0.277 0.835 0.656 0.013 0.370 0.767 0.562
FLSW-100 0.000 0.500 0.667 0.433 0.000 0.497 0.670 0.411
FLSW-200 0.000 0.410 0.742 0.452 0.003 0.557 0.613 0.360
FLSW-300 0.003 0.573 0.597 0.342 0.003 0.827 0.295 0.254
FLSW-400 0.003 0.810 0.318 0.244 0.007 0.943 0.107 0.195
WaterSeeker(Ours) 0.007 0.287 0.829 0.642 0.010 0.390 0.753 0.542
In Table 1, we present the detection performance of WaterSeeker alongside various baseline algorithms, evaluated on four datasets. Given that full-text detection lacks localization capabilities, its FNR is necessarily 1 under the evaluation rule requiring IoU $>\theta$ for successful detection. Therefore, we additionally report the FNR and corresponding F1 score for Full-text Detection without the IoU $>\theta$ constraint, highlighted in red.
The results demonstrate that WaterSeeker achieves lower FNR and higher F1 score, significantly surpassing other watermark detection algorithms, including Full-text Detection, FLSW-100, FLSW-200, FLSW-300, and FLSW-400, while being comparable to WinMax. Based on the principle that WinMax is guaranteed to reach the gold index by evaluating all possible windows, it represents the upper bound of detection performance. However, subsequent experiments show that WinMax has extremely high time complexity (Section 6.4). Furthermore, the IoU [^1] For samples with no detected segments, the IoU is 0. For samples with multiple detected segments, take the max(IoU). results in the table reveal that WaterSeeker possesses a significantly higher localization capability compared to the baselines, again comparable to WinMax.
6.3 Adaptability Analysis
While the four datasets used in the main experiment incorporated a mix of watermarked segments with varying strengths and lengths, this section presents a more granular comparison of the detection capabilities of WaterSeeker and the FLSW method (FLSW-100, FLSW-200, FLSW-300, and FLSW-400) across specific watermark strengths and lengths. As shown in Figure 3, the performance of different detection algorithms exhibits a consistent trend: watermarked segments with higher strengths and longer lengths are more easily detected. Moreover, it is evident that FLSW performs well within a length range close to its window size, but shows poor adaptability in other ranges. This limitation arises from its fixed-length nature, which leads to dilution effects when detecting longer or shorter watermarked segments, thereby reducing detection performance. In contrast, WaterSeeker exhibits good adaptability across different length ranges by using anomaly extraction techniques. Additionally, it incorporates a top-k score mechanism to adapt to varying watermark strengths.
(a) KGW, $\delta=2.0$
(b) KGW, $\delta=1.5$
(c) KGW, $\delta=1.0$
(d) Aar, temp=0.3
(e) Aar, temp=0.2
(f) Aar, temp=0.1
Figure 3: This figure compares the detection performance of WaterSeeker and FLSW under varying watermark lengths and strengths, using Llama-2-7b as the generation model.
6.4 Time Complexity Analysis
Table 2: Average detection time per sample for various detection methods ($T\approx 10,000$, Unit: s).
Full-text Detection WinMax FLSW-100 FLSW-200 FLSW-300 FLSW-400 WaterSeeker
KGW 1.70 14.16 1.76 1.76 1.79 1.80 1.75
Aar 0.54 2733.78 1.62 1.65 1.66 1.64 1.68
Table 2 lists the average time spent per sample for WaterSeeker and other baseline algorithms during detection. The results show that the time taken by WaterSeeker is comparable to that of full-text detection and FLSW, and significantly lower than the time required by WinMax. Sections 4.2 and 5.4 analyze the theoretical time complexity of WinMax and WaterSeeker, respectively. This section aims to offer a more comprehensive analysis based on experimental results, taking into account the constant factors preceding the complexity terms. Additionally, it will investigate the reasons behind the significant disparity in execution time of WinMax between KGW and Aar.
In this experimental setting, let the time required to compute a score in the score list be $t_{1}$, and the time required to compute the watermark detection statistic within a local window be $t_{2}$. The time expenditure for the WinMax [^2] The WinMax algorithm, in practical implementation, also requires computing the score list for the entire text first, then traversing the windows to calculate the detection statistic. algorithm to detect a sample can be expressed as $N\cdot t_{1}+N\cdot(W_{max}-W_{min})\cdot t_{2}$. For WaterSeeker, the first step of localization consists of three sub-steps: score list computation, anomaly detection, and fragment connection. Each of these sub-steps has a complexity of $O(N)$. If we denote the time required for the latter two sub-steps as $t_{3}\cdot N$, then the total time for the first step is $N\cdot t_{1}+N\cdot t_{3}$. The second step, local traverse detection, requires $W^{2}\cdot t_{2}$ time. Consequently, the total time expenditure for the entire process can be expressed as $N\cdot t_{1}+N\cdot t_{3}+W^{2}\cdot t_{2}$.
For the Aar (Aaronson & Kirchner, 2022) watermarking method, calculating the watermark detection statistic involves computing a GammaTransform function, which is a relatively complex computation, resulting in a larger value for $t_{2}$. In this case, the time difference between WinMax and WaterSeeker will become significant due to the large disparity in the coefficients preceding $t_{2}$. In this experiment, with $T=10,000,W_{max}=400,W_{min}=100,W^{2}=2,500$, the coefficient difference amounts to a factor of 1,200. For the KGW (Kirchenbauer et al., 2023) watermarking method, calculating the z-score within a local window is a relatively straightforward computation. In this case, the time difference between WinMax and WaterSeeker will be reduced.
Through the above analysis, it is clear that WaterSeeker is a general watermark detection method with consistently low time expenditure across various watermarking algorithms. In contrast, WinMax exhibits extremely high time costs in situations where calculating the watermark detection statistic is complex, making it challenging for practical use.
6.5 Ablation Study
Table 3: This table presents the contributions of the first step of WaterSeeker: Suspicious Segment Localization. It lists the average coverage of localization results relative to the ground truth segments under various watermark algorithms and watermark strengths, as well as the average offsets of the detected start and end indices.
Metrics KGW Aar
$\delta=2.0$ $\delta=1.5$ $\delta=1.0$ temp=0.3 temp=0.2 temp=0.1
Average Coverage 0.992 0.984 0.953 0.996 0.991 0.977
Average Offset (Start) 20.799 30.057 13.992 22.559 30.546 20.624
Average Offset (End) 25.985 25.299 28.030 29.459 33.059 30.304
In this section, we will analyze the effectiveness of the two stages of WaterSeeker through an ablation study. The first stage of WaterSeeker, Suspicious Segment Localization, aims to achieve significant coverage of the ground truth segments while keeping the start and end offsets within a specified window size. This ensures that subsequent local traversals can reach the gold index. Table 3 illustrates the performance of Step 1, showing that the localization results of Suspicious Segment Localization achieve an average coverage level exceeding 0.95 across various watermark algorithms and strengths. Additionally, both the average start offset and average end offset are less than 50, remaining within the designated window size. This indicates that Step 1 fulfills its intended purpose effectively.
Local Traverse Detection performs a localized traversal based on the segments narrowed down in the first step, allowing for more refined verification within the window. Table 4 illustrates the differences in detection outcomes when using Local Traverse compared to directly applying full-text detection with the localization results. It can be found that across different LLMs and watermarking algorithms, Local Traverse consistently enhances the detection performance, making it an indispensable component of WaterSeeker.
Table 4: This table compares the detection performance of using Local Traverse versus not using it across the four datasets.
Model Setting KGW Aar
FPR FNR F1 IoU FPR FNR F1 IoU
Llama-2-7b w. local traverse 0.017 0.303 0.813 0.629 0.010 0.433 0.719 0.510
w/o local traverse 0.003 0.390 0.756 0.520 0.000 0.450 0.710 0.462
Mistral-7b w. local traverse 0.007 0.287 0.829 0.642 0.010 0.390 0.753 0.542
w/o local traverse 0.003 0.380 0.763 0.534 0.003 0.413 0.738 0.490
7 Further Application
Figure 4: An illustration of WaterSeeker applied in AI detection system.
WaterSeeker not only provides accurate detection results, but its localization capabilities also contribute to building a more transparent and interpretable AI detection system. As illustrated in Figure 4, the AI detection system powered by WaterSeeker receives documents for analysis and outputs three components: (1) Detection Result, which indicates whether AI assistance was utilized; (2) Suspicious AI-generated Segments, which highlights segments identified as potentially AI-generated; and (3) AI Ratio, which displays the proportion of content attributed to AI. For example, in the context of academic integrity, highlighting suspicious segments provides actionable evidence for assessments and appeals. Additionally, reporting the AI ratio allows for more flexible establishment of academic misconduct standards.
8 Conclusion
In this work, we defined a new scenario for detecting watermarked segments in large documents and established its evaluation metrics. We identified the limitations of traditional full-text detection methods in this context and proposed a “first locate, then detect” watermark detection algorithm that utilizes a coarse-to-fine strategy. We validated the detection performance and time complexity of our algorithm through a series of analyses and experiments, demonstrating its ability to effectively balance both aspects. For future improvements, we recommend exploring more advanced locating methods based on this concept, which may yield better detection results.
Appendix A Pseudocode of Detection Baselines
Algorithm 2 WinMax Algorithm
1:procedure WinMaxDetection(tokens, $W_{min}$, $W_{max}$, threshold)
2:     hasWatermark $\leftarrow$ False, indices $\leftarrow$ [ ]
3:     maxStat $\leftarrow$ - $\infty$, bestIndex $\leftarrow$ None
4:     for $W$in [$W_{min}$, $W_{max}$do
[…]
6:              stat $\leftarrow$ CalculateStatistics(tokens[$i:i+W$])
7:              if stat $>$ maxStat then
8:                  maxStat $\leftarrow$ stat
9:                  bestIndex $\leftarrow$ $(i,i+W)$
10:              end if
11:         end for
12:     end for
13:     if maxStat $>$ threshold then
14:         hasWatermark $\leftarrow$ True
15:         indices.append(bestIndex)
16:     end if
17:     return hasWatermark, indices
18:end procedure
Algorithm 3 FLSW Algorithm
1:procedure FLSWDetection(tokens, $W$, threshold)
2:     hasWatermark $\leftarrow$ False
3:     indices $\leftarrow$ [ ]
[…]
5:         stat $\leftarrow$ CalculateStatistics(tokens[$i:i+W$])
6:         if stat $>$ threshold then
7:              hasWatermark $\leftarrow$ True
8:              indices.append(($i$, $i+W$))
9:         end if
10:     end for
11:     indices $\leftarrow$ ConnectFragments(indices)
12:     return hasWatermark, indices
13:end procedure
Appendix B Detailed Proof for Aar
B.1 Expected p-value of watermarked text
In this sub-section, we aim to analyze how the expected p-value for watermarked text varies with changes in the detection window size $W$.
Recall the p-value calculation formula:
[…]
where $S=\sum_{i=1}^{W}\log(\frac{1}{1-u_{i}})$, and $W$ is the window size.
For watermarked tokens, $E[\log(\frac{1}{1-u_{i}})]=\mu_{1}$ , and for non-watermarked tokens, $E[\log(\frac{1}{1-u_{i}})]=\mu_{0}$, where $\mu_{1}>\mu_{0}$.
The expectation of S for different W values is as follows:
  • When $W\leq L$: $E[S]=W\mu_{1}$
  • When $W>L$: $E[S]=L\mu_{1}+(W-L)\mu_{0}$
The expectation of p-value is:
$E[\text{p-value}]=E[\text{GammaTransform}(S,W,\text{loc}=0,\text{scale}=1)]$ (9)
Since GammaTransform is non-linear, we cannot directly substitute $E[S]$. However, we can use Jensen’s inequality to obtain an approximation:
$E[\text{GammaTransform}(S,W,0,1)]\geq\text{GammaTransform}(E[S],W,0,1)$ (10)
Therefore, we can analyze $\text{GammaTransform}(E[S],W,0,1)$ to obtain a lower bound for the expectation of p-value. Define function $f(W)=\text{GammaTransform}(E[S],W,0,1)$:
  • When $W\leq L$: $f(W)=\text{GammaTransform}(W\mu_{1},W,0,1)$
  • When $W>L$: $f(W)=\text{GammaTransform}(L\mu_{1}+(W-L)\mu_{0},W,0,1)$
Analyzing the behavior of $f(W)$:
  • When $W\leq L$, both $E[S]$ and the shape parameter W increase as $W$ increases.
  • When $W$ just exceeds $L$, the growth rate of $E[S]$ suddenly decreases (from $\mu_{1}$ to $\mu_{0}$), while the shape parameter W continues to increase linearly.
Consider the behavior of $f(W)$ near $W=L$:
  • When $W$ increases from $L-\epsilon$ to $L$, $E[S]$ increases by $\epsilon\mu_{1}$, and the shape parameter increases by $\epsilon$.
  • When $W$ increases from $L$ to $L+\epsilon$, $E[S]$ increases by $\epsilon\mu_{0}$, and the shape parameter increases by $\epsilon$.
Since $\mu_{1}>\mu_{0}$, at the point $W=L$, the rate of decrease of $f(W)$ suddenly slows down. This suggests that $f(W)$ is likely to reach its minimum value at $W=L$, or at a point very close to $L$. While this analysis does not strictly prove that the expectation of p-value is minimized exactly at $W=L$, it strongly suggests that the expectation of p-value is likely to reach its minimum value at or very near $W=L$ [^3] For a more rigorous proof, a deeper mathematical analysis of the GammaTransform function or numerical simulations would be necessary.
B.2 Calculation of p-threshold to control false positive rate
Similar to the analysis for KGW method, we also need to analyze the constraints on the p-value threshold $p^{*}$ when the false positive rate within the specified window is set to be lower than a target value $\alpha$.
For non-watermarked text, $u_{i}\sim\text{Uniform}([0,1])$. Consequently, the test statistic $S$ follows a Gamma distribution: $S\sim\text{Gamma}(W,1)$, where $W$ is the shape parameter and 1 is the scale parameter. The p-value is calculated using the Gamma CDF:
[…]
where GammaCDF is the cumulative distribution function of the Gamma distribution with shape parameter $W$ and scale parameter 1. To achieve a false positive rate of $\alpha$, we need to set a threshold $p^{*}$ such that: $P(\text{p-value}<p^{*})=\alpha$.
Given the definition of p-value, this is equivalent to: $P(1-\text{GammaCDF}(S,W,1)<p^{*})=\alpha$, which can be rewritten as: $P(S>\text{GammaInv}(1-p^{*},W,1))=\alpha$, where GammaInv is the inverse of the Gamma CDF.
Since $S$ follows a $\text{Gamma}(W,1)$ distribution for non-watermarked text, we can express this as:
$1-\text{GammaCDF}(\text{GammaInv}(1-p^{*},W,1),W,1)=\alpha.$ (12)
Solving this equation for $p^{*}$, we get $p^{*}=\alpha$, which is also a constant value for different $W$.
Appendix C Detail of Threshold Selection Within the Specified Window
A key role of threshold selection is to control the false positive rate. In this context, the task involves detecting watermark fragments within long texts, which requires traversing extensive content and can lead to an accumulation of false positives. Therefore, managing the false positive rate within the detection window is crucial in this scenario. In the experiment, we set the targeted false positive rate $\alpha$ within the detection window to $10^{-6}$.
C.1 Rationale for Setting $\alpha$ to $10^{-6}$
Table 5: Simulated FPR of WaterSeeker using 10,000 samples for each watermarking method. The targeted false positive rate within the detection window is set to $10^{-6}$.
Watermarking Method Simulated FPR
KGW 0.0054
Aar 0.0042
WaterSeeker, WinMax, and FLSW all involve employing sliding windows for text traversal and conduct full-text detection within each window. As these windows overlap, they cannot be treated as independent, making it challenging to derive a theoretical upper bound for the document-level FPR from the targeted FPR within each window. Given this, we utilize large-scale data simulation to demonstrate that, with a targeted false positive rate of $10^{-6}$ within each window, our proposed method WaterSeeker maintains an acceptable false positive rate.
For the KGW method, we set $\gamma=0.5$ in our experiments, meaning each token in non-watermarked text has a 0.5 probability of being green and 0.5 probability of being red. In the simulation, we generate 10,000 samples, each containing 10,000 tokens, with each token having a 0.5 probability of being 1 and 0.5 probability of being 0. For the Aar method, each token in non-watermarked text corresponds to $u_{i}\sim\text{Uniform}[0,1]$. In the simulation, we again generate 10,000 samples, each containing 10,000 tokens, with each token randomly assigned a floating-point number from $[0,1]$.
We then apply WaterSeeker to detect watermarked segments within these samples, setting the targeted false positive rate within the detection window to $10^{-6}$. The large-scale simulation results in Table 5 demonstrate that WaterSeeker maintains a false positive rate of approximately 0.005, which is considered acceptable. For scenarios requiring more stringent FPR control, the targeted false positive rate can be adjusted downward. However, this inevitably compromises the detection rate, highlighting a key challenge in watermarked segment detection within large documents.
C.2 Setting the Threshold to Achieve a Target False Positive Rate $\alpha$
For KGW, as analyzed in Section 5.1, when the window size is large, we can approximate using the Central Limit Theorem, resulting in $z^{*}=\Phi^{-1}(1-\alpha)$. When $\alpha=10^{-6}$, this gives $z\approx 4.75$. However, when the window size $W$ is small, the approximation to a normal distribution using the Central Limit Theorem may lead to significant deviations. Therefore, we will use the binomial distribution for precise calculations. $x\sim B(W,\gamma)$ describes the number of green tokens in a window of size $W$ follows a binomial distribution, therefore:
$z=\frac{x-\gamma W}{\sqrt{W\gamma(1-\gamma)}}.$
To find $P(z\geq z^{*})$:
$P(z\geq z^{*})=P\left(\frac{x-\gamma W}{\sqrt{W\gamma(1-\gamma)}}\geq z^{*}\right).$
Expanding this, we have:
$P(z\geq z^{*})=\sum_{k=0}^{W}\binom{W}{k}\gamma^{k}(1-\gamma)^{W-k}\mathbb{I}\left\{\frac{k-\gamma W}{\sqrt{W\gamma(1-\gamma)}}\geq z^{*}\right\}.$
This is the exact expression for $P(z\geq z^{*})$ without any approximations.
We can further simplify:
$P(z\geq z^{*})=\sum_{k=0}^{W}\binom{W}{k}\gamma^{k}(1-\gamma)^{W-k}\mathbb{I}\left\{k\geq\gamma W+z^{*}\sqrt{W(1-\gamma)}\right\}.$
We need to find an appropriate $z^{*}$ such that $P(z\geq z^{*})<\alpha$. This function does not have a direct analytical solution, so we can increment $z^{*}$ in steps of 0.01 until the probability exceeds $\alpha$. The final value of $z^{*}$ is dependent on $W$, and we pre-compute these values during experiments and store them in a dictionary. In experiments, for detected segments with a length of 200 or more, we directly apply the Central Limit Theorem approximation, setting $z=4.75$. For segments shorter than 200, we use the binomial distribution and retrieve the corresponding threshold from the pre-computed dictionary.
For Aar, as analyzed in Appendix B, $p^{*}=\alpha$, therefore set $p^{*}=10^{-6}$ for all lengths.
Appendix D Sample Distribution in the Main Experiment
It can be observed from Figure 5 that all four datasets include samples of watermarked segments with varying intensities and lengths, demonstrating the comprehensiveness and fairness of the dataset construction.
(a) KGW, Llama
(b) Aar, Llama
(c) KGW, Mistral
(d) Aar, Mistral
Figure 5: Distribution of sample numbers for the four datasets involved in the main experiment.